Evaluate b to the power of 2-2bc+c to the power of 2. If b=3 and c=-2.
Solve 10x-4-5x=6+3x-8

Thanks
(two different problems)

Answer by rajagopalan(174)

Problem 1: b to the power of 2-2bc+c to the power of 2. If b=3 and c=-2.
2-2bc+c=2-(-12)-2=14-2=12
power2=24
3 power 12 power 2
=3^24
*************
*************
Problem 2
10x-4-5x=6+3x-8
10x-5x-3x=6+4+8
2x=2
x=1
ok